# Control Your Robot: A Unified System for Robot Control and Policy Deployment

### v1

@@ I Introduction
Recent progress in data-driven embodied-intelligence models has markedly advanced robotic control [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], establishing embodied intelligence as a central research paradigm. These gains are enabled by the availability of large-scale robot-manipulation datasets and the adoption of foundation models. Representative methods, including ACT [11], G3Flow [12], Diffusion Policy [13], ManiCM [14], OpenVLA [15], RDT [16], and $\pi_{0}$ [17], learn robust perceptual and control representations from human demonstrations and diverse corpora, and consequently achieve strong generalization across tasks, objects, and embodiments.
@@
Control Your Robot is released as a fully open-source framework, providing APIs for data collection as well as training and deployment code. Control Your Robot available at https://control-your-robot.git.
@@ II-A Robot Data Collection System
Teleoperation remains a practical approach for collecting demonstrations, particularly in contact-rich tasks. SharedAssembly [18] enables efficient tele-assembly, while PATO [19] employs policy assistance to reduce operator effort and scale across robots. Platforms like RAPID Hand [20] integrate visual, proprioceptive, and tactile sensing for fine-grained manipulation. Building on these principles, our framework provides a modular, low-latency pipeline for scalable demonstration capture and training.
@@ II-B Robot Data Storage System
Efficient storage is critical for reproducibility and scalability. LeRobot [21] offers a unified format with high-throughput access, while large-scale efforts like RT-X [22] enable foundation-level learning over millions of trajectories. Such frameworks address indexing, compression, and retrieval. Our system complements these by providing a streamlined interface between collection and long-term management.
@@ II-C General Embodied Model Deployment
Deploying embodied models requires real-time perception–action loops and robust integration with dynamic environments. Existing frameworks typically separate inference, sensing, and control, often leveraging middleware like ROS [23]. Standardized interfaces and modular designs support policy transfer across robots and tasks [24, 25]. Recent systems, such as RAI [24] and ROS-LLM [26], extend this paradigm by integrating large language models and structured reasoning for intuitive task programming. Similarly, ROS-X-Habitat [25] bridges ROS and Habitat for seamless simulation–real transfer. Building on these ideas, our system offers a modular deployment interface aligned with data collection and storage, enabling reproducible real-world evaluation.
@@ III-C3 Process Standardization
Figure 3: Evaluated Tasks. We assess distinct manipulation tasks: Place Can on Pot (place a can on a pot edge); Pick Dual Bottles (simultaneously grasp two irregular bottles); Put Cup in Cabinet (handle recognition, drawer operation, placement, and closure); Stack Two Bowls (sequential pick-and-stack).
@@
  • Can the data collected via Control Your Robot maximally enhance the performance of end-to-end embodied intelligence policies?
@@ IV-A Implementation Details and Task Description
We conducted experiments on real-world robotic platforms, including both single-arm and dual-arm systems, across four representative manipulation tasks (see Fig. 3 for real-world demonstrations). The single-arm robotic system consists of a slave Dr ALOHA arm and a master AigleX piper arm. The slave arm is equipped with a wrist-mounted camera and a global camera, together forming a single-arm teleoperation system. The dual-arm robotic system consists of two Y1 ALOHA arms configured in a master–slave integrated manner. Each arm is equipped with a wrist-mounted camera, and a shared global camera is positioned between the two arms, together forming a tabletop dual-arm teleoperation system.
All sensor data and robot state information were collected at 30 Hz. For policy learning, we employed ACT (a widely used imitation learning algorithm) and PI (an advanced vision-language-action model). For ACT training, we used the AdamW optimizer for 60k iterations. For PI fine-tuning, we applied a cosine learning rate scheduler for 30k iterations. Unless otherwise specified, all experiments were evaluated over 50 test trials for success rate. Our method was assessed on a diverse set of tasks, each designed to test distinct robotic skill sets:
@@
  • Stack Bowls Three: A high-precision, long-horizon manipulation task that requires the model to accurately stack three bowls in the center of the workspace, following a strict order: right before left, and near before far.
@@ IV-B Low-Latency Data Collection
We evaluated serial and parallel acquisition under a 60 Hz requirement. The effective frame rates were 59.28 Hz (serial) and 59.95 Hz (parallel), showing that parallel acquisition mitigates delay from mixed-frequency devices and ensures real-time data collection.
@@ IV-C Model Deployment and Data Replay
For deployment, 50 trajectories were executed and compared with expert demonstrations, showing high consistency with low variance and mean deviation. In data replay, ACT and PI models were trained on 20 sampled sequences and accurately reproduced expert strategies (Fig. LABEL:fig:data). These results confirm that data from Control Your Robot effectively supports fine-grained policy learning.

### v2

@@ I Introduction
Recent progress in data-driven embodied-intelligence models has markedly advanced robotic control [kawaharazuka2025vision, long2025survey, mu2025robotwin, chen2025robotwin, chen2025benchmarking, liu2025avr, fu2025cordvip, motoda2025recipe], establishing embodied intelligence as a central research paradigm. These gains are enabled by the availability of large-scale robot-manipulation datasets and the adoption of foundation models. Representative methods, including ACT [fu2024mobile], G3Flow [chen2025g3flow], Diffusion Policy [chi2023diffusion], ManiCM [lu2024manicm], OpenVLA [kim2024openvla], RDT [liu2024rdt], and $\pi_{0}$ [black2024pi_0], learn robust perceptual and control representations from human demonstrations and diverse corpora, and consequently achieve strong generalization across tasks, objects, and embodiments.
@@
Control Your Robot is released as a fully open-source framework, providing APIs for data collection as well as training and deployment code. Control Your Robot available at https://control-your-robot.
@@ II-A Robot Data Collection System
Teleoperation remains a practical approach for collecting demonstrations, particularly in contact-rich tasks. SharedAssembly [sharedassembly2025] enables efficient tele-assembly, while PATO [pato2023] employs policy assistance to reduce operator effort and scale across robots. Platforms like RAPID Hand [rapidhand2025] integrate visual, proprioceptive, and tactile sensing for fine-grained manipulation. Building on these principles, our framework provides a modular, low-latency pipeline for scalable demonstration capture and training.
@@ II-B Robot Data Storage System
Efficient storage is critical for reproducibility and scalability. LeRobot [lerobot2024] offers a unified format with high-throughput access, while large-scale efforts like RT-X [rtx2023] enable foundation-level learning over millions of trajectories. Such frameworks address indexing, compression, and retrieval. Our system complements these by providing a streamlined interface between collection and long-term management.
@@ II-C General Embodied Model Deployment
Deploying embodied models requires real-time perception–action loops and robust integration with dynamic environments. Existing frameworks typically separate inference, sensing, and control, often leveraging middleware like ROS [ros-hub]. Standardized interfaces and modular designs support policy transfer across robots and tasks [rai2025, ros-x-habitat2021]. Recent systems, such as RAI [rai2025] and ROS-LLM [ros-llm2024], extend this paradigm by integrating large language models and structured reasoning for intuitive task programming. Similarly, ROS-X-Habitat [ros-x-habitat2021] bridges ROS and Habitat for seamless simulation–real transfer. Building on these ideas, our system offers a modular deployment interface aligned with data collection and storage, enabling reproducible real-world evaluation.
@@ III-C3 Process Standardization
  • How does ”Control Your Robot” systematically contribute to policy robustness through enhanced data reliability?
@@ IV-A Implementation Details and Task Description
We conducted experiments on real-world robotic platforms, including both single-arm and dual-arm systems, across four representative manipulation tasks (see Fig. LABEL:fig:data for real-world demonstrations). The single-arm robotic system consists of a slave DR-ALOHA arm and a master AGILEX-PIPER-ALOHA The slave arm is equipped with a wrist-mounted camera and a global camera, together forming a single-arm teleoperation system. The dual-arm robotic system consists of two IMETA-Y1-ALOHA arms configured in a master–slave integrated manner. Each arm is equipped with a wrist-mounted camera, and a shared global camera is positioned between the two arms, together forming a tabletop dual-arm teleoperation system.
All sensor data and robot state information were collected at 30 Hz. For policy learning, we employed ACT (a widely used imitation learning algorithm) and PI0 (an advanced vision-language-action model). For ACT training, we used the AdamW optimizer for 60k iterations. For PI fine-tuning, we applied a cosine learning rate scheduler for 30k iterations. Unless otherwise specified, all experiments were evaluated over 50 test trials for success rate. Our method was assessed on a diverse set of tasks, each designed to test distinct robotic skill sets:
  • Place Can Pot: A spatial coordination task where the robot moves a sauce can from the table to a precise position next to the pot, requiring accurate placement and smooth motion control.
  • Pick Dual Bottles: A dual-arm coordination task where each arm grasps a separate drink bottle simultaneously, emphasizing balance, synchronization, and grasp stability.
@@
  • Stack Two Bowls: A high-precision, long-horizon manipulation task that requires the model to accurately stack three bowls in the center of the workspace, following a strict order: right before left, and near before far.
@@ IV-B Low-Latency Data Collection
We additionally tested data acquisition with multiple sampling frequencies: the robotic arm was sampled at 600Hz, and images at 60Hz, using a dual-arm, three-view setup. The resulting effective sampling rates were 599.247Hz for the robotic arm and 59.997Hz for images, demonstrating that our system can reliably handle mixed-frequency data collection.
We also evaluated a multi-frequency control setup that combines high-frequency teleoperation with low-frequency data collection. Specifically, we used 300 Hz control for master–slave arm synchronization and 60 Hz for recording robot states and real-time images (the latter limited by the camera’s maximum frame rate). Over 10 continuous runs, the average real synchronization frequency stabilized at 282.35 Hz, and the data collection frequency at 59.99 Hz.
@@ IV-C Model Deployment and Data Replay
During the deployment testing phase, 50 operational trajectories were executed and quantitatively compared with expert demonstration data. The results indicate a high degree of consistency between the generated trajectories and the expert data, with both variance and mean deviation maintained at low levels. In the data replay phase, both the ACT and PI models were trained using 20 sampled sequences (note: the PI model was excluded from graphical comparisons as it was trained on only a single task). Both models successfully replicated the expert operational strategies, as shown in (Fig. LABEL:fig:data). With the exception of the Put Cup Cabinet task, the actual robot trajectories for the remaining tasks exhibited significant stochasticity, thereby imposing greater demands on the models’ generalization capability. Owing to its limited parameter size, the ACT model demonstrated insufficient learning performance on real robot data, leading to notable performance deviations. Overall, the experimental findings confirm that the data collected by the ”Control Your Robot” system effectively supports fine-grained policy learning.